OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 16? x2 + y2 = 4 x2 + y2 = 16 x2 + y2 = 256

OpenStudy (hoblos):

x2 + y2 =16^2 x2 + y2 = 256